OpenStudy (anonymous):

how to find integration of ln(x) with respect to x, approximation is allowed.

OpenStudy (anonymous):

graphical or any kind of logical solution is acceptable....

OpenStudy (anonymous):

Use integration by parts.. \[\int\limits \ln(x)dx \rightarrow f=x, F=x, g=\ln(x), g'=\frac{1}{x}\] \[xln(x)-\int\limits x \frac{1}{x}dx \rightarrow xln(x)-\int\limits dx \rightarrow xln(x)-x+C \rightarrow x[\ln(x)-1]+C\]

OpenStudy (anonymous):

thank u very much...

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

but if it is a definite integral problem and x->0. then how would you approach?

OpenStudy (anonymous):

sorry made a typo litte f=1

OpenStudy (anonymous):

read this link about improper integrals.... it also has examples http://tutorial.math.lamar.edu/Classes/CalcII/ImproperIntegrals.aspx

OpenStudy (anonymous):

as far i understood it should go to infinity..